### v2

# Learning Representations of Event Time Series with Sparse Autoencoders for Anomaly Detection, Similarity Search, and Unsupervised Classification

Abstract
Event time series are sequences of discrete events occurring at irregular time intervals, each associated with a domain-specific observational modality. They are common in domains such as high-energy astrophysics, computational social science, cybersecurity, finance, healthcare, neuroscience, and seismology. Their unstructured and irregular structure poses significant challenges for extracting meaningful patterns and identifying salient phenomena using conventional techniques. We propose novel two- and three-dimensional tensor representations for event time series, coupled with sparse autoencoders that learn physically meaningful latent representations. These embeddings support a variety of downstream tasks, including anomaly detection, similarity-based retrieval, semantic clustering, and unsupervised classification. We demonstrate our approach on a real-world dataset from X-ray astronomy, showing that these representations successfully capture temporal and spectral signatures and isolate diverse classes of X-ray transients. Our framework offers a flexible, scalable, and generalizable solution for analyzing complex, irregular event time series across scientific and industrial domains.
Machine Learning, ICML
1 Introduction
Event time series—irregular sequences of timestamped discrete events of a particular observational modality—arise across a wide range of scientific and industrial domains. In astronomy, they represent photon arrivals detected by telescopes; in computational social science, they trace social media activity or user interactions; in cybersecurity, they log system alerts; in finance, they capture transactions and market fluctuations; in healthcare, they reflect patient monitoring data; in neuroscience, they record neuronal spike trains; and in seismology, they track seismic wave arrivals. These time series encode rich temporal and domain-specific information, often containing salient patterns, natural groupings, and rare anomalies. Yet this information can remain inaccessible, as the irregular and unstructured nature of event time series limits the effectiveness of conventional analysis techniques. Unlocking their latent structure requires more flexible and expressive methods that simultaneously enable a range of follow-on analyses. In this work, we present a representation learning (Bengio et al., 2013) framework based on sparse autoencoders (SAEs; Hinton & Salakhutdinov 2006; Ng et al. 2011) to extract meaningful features from event time series, enabling downstream tasks such as anomaly detection, similarity searches, clustering, and unsupervised classification. By enforcing sparsity in the latent space, the model is encouraged to focus on the most physically relevant features, while remaining robust to nuisance variation arising from noise, observational systematics, or contextual artifacts unrelated to the underlying phenomena.
We demonstrate our pipeline on archival data from the Chandra X-ray Observatory (Weisskopf et al., 2000). In X-ray astronomy, event files record the arrival times and energies of individual photons, capturing the stochastic behavior of astrophysical sources. Some of the most interesting discoveries—such as extragalactic fast X-ray transients (FXTs; Jonker et al. 2013; Glennie et al. 2015; Xue et al. 2019; Lin et al. 2022) and even the first candidate extragalactic planet (Di Stefano et al., 2021)—have been made serendipitously, often years after the data were collected. FXTs are brief, intense X-ray flares from extragalactic sources lasting only minutes to hours, whereas the planet candidate was identified as a dip in the X-ray light curve. These and other transient sources span a wide range of timescales and spectral properties, offering insights into their astrophysical origins. However, conventional methods often rely on rigid assumptions and handcrafted summary statistics (Yang et al., 2019; Quirola-Vásquez et al., 2022), limiting their ability to capture the full complexity of such phenomena. In contrast, our sparse representation learning approach provides a flexible and scalable alternative—well-suited for anomaly detection, transient searches, and unsupervised classification in large datasets containing diverse types of event time series. Our method has already led to the discovery of XRT 200515 (Dillmann et al., 2025), a previously unknown FXT and a new hyperluminous supersoft X-ray source (Sacchi et al., 2025) that had been buried in the archive for years.
2 Method
2.1 Tensor Representation of Event Time Series
For a dataset with $n$ event time series $\{\boldsymbol{x}_{i}\}^{n}_{i=1}$ that vary in length $N$ and duration $T$, we introduce a fixed-size tensor representation that standardizes the data for subsequent analysis. This approach is inspired by the DMDT maps developed for optical light curves by Mahabal et al. (2017).
Assume that each event time series consists of a sequence of timestamps $\boldsymbol{t}=\{t_{k}\}_{k=1}^{N}$ and associated event modality values $\boldsymbol{E}=\{E_{k}\}_{k=1}^{N}$. In the context of X-ray astronomy, $\boldsymbol{t}$ would correspond to the photon arrival times and $\boldsymbol{E}$ to the photon energies. The total duration of the event time series is given by $T=t_{N}-t_{1}$, and the event values lie in the range $E_{\min},E_{\max}]$, where $E_{\min}=\min_{k}E_{k}$ and $E_{\max}=\max_{k}E_{k}$. To standardize the inputs across event time series, we normalize the time axis as
$\boldsymbol{\tau}=\frac{\boldsymbol{t}-t_{1}}{T},$ (1)
which maps all timestamps to the normalized interval $\boldsymbol{\tau}\in[0,1]$. For the event values, we apply a transformation
$\boldsymbol{\epsilon}=f(\boldsymbol{E}),$ (2)
with $\boldsymbol{\epsilon}\in[\epsilon_{\min},\epsilon_{\max}]$, where $\epsilon_{\min}=f(E_{\min})$ and $\epsilon_{\max}=f(E_{\max})$. The choice of transformation function $f(\cdot)$ depends on the domain and distribution of the event modality and is intended to highlight relevant structure for downstream analysis. To capture local temporal dynamics beyond absolute event times, we introduce a third dimension based on inter-event intervals. Specifically, we compute the time differences between consecutive events, $\boldsymbol{\Delta t}=\{\Delta t_{k}\}_{k=1}^{N}=\{t_{k+1}-t_{k}\}_{k=1}^{N-1}$, and normalize them to the range $[0,1]$ as
$\boldsymbol{\Delta\tau}=\frac{\boldsymbol{\Delta t}-\Delta t_{\min}}{\Delta t_{\max}-\Delta t_{\min}},$ (3)
where $\Delta t_{\min}=\min_{k}\Delta t_{k}$ and $\Delta t_{\max}=\max_{k}t_{k}$. This third axis captures local temporal density and is a proxy for the event rate: small values of $\Delta\tau$ correspond to periods of concentrated activity, while larger values indicate sparser or quiescent intervals. In heterogeneous, irregular discrete time series, temporal spacing often conveys essential structural information beyond the absolute timestamps alone. Including this dimension improves the expressiveness of the representation, allowing models to capture transient dynamics and distinguish between different temporal regimes.
These three normalized axes—event time ($\boldsymbol{\tau}$), transformed modality ($\boldsymbol{\epsilon}$), and inter-event interval ($\boldsymbol{\Delta\tau}$)—are discretized into fixed-width bins with $n_{\tau}$, $n_{\epsilon}$, and $n_{\Delta\tau}$ bins, respectively. Each event is assigned to a voxel in the resulting three-dimensional histogram based on its coordinates in this normalized space. The number of bins along each axis can be chosen flexibly to achieve the desired level of temporal or modality-specific resolution, depending on the application, domain, and characteristics of the data. This allows the representation to be tailored to the appropriate scale and granularity for a given task, while still providing a consistent input format for models, even when the dataset contains sequences of variable length and irregular temporal structure. This results in a tensor $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{\tau}\times n_{\epsilon}\times n_{\Delta\tau}}$, which provides a standardized, structured input format for each of the original event time series $\boldsymbol{x}_{i}$. Each tensor is then collected into a final processed dataset $\{\boldsymbol{X}_{i}\}^{n}_{i=1}$. We also refer to these tensors as EEttdtdt cubes. If we were to only bin along time $\boldsymbol{\tau}$ and modality $\boldsymbol{\epsilon}$, we would obtain two-dimensional tensors $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{\tau}\times n_{\epsilon}}$, which we refer to as EEtt maps. Appendix A shows EEtt maps and EEttdtdt cubes for an X-ray flare, dip and pulsating source.
Figure 1: Two-dimensional t-SNE projection of the learned latent space from the SAE applied on the EEtt-dtdt cubes. Panel A: Points are color-coded by the variability index of the corresponding X-ray sources. Panel B: Points are color-coded by the hard-to-soft X-ray hardness ratio. Panel C: Known dips, flares, and pulsating sources (crosses) occupy distinct clusters in the embedding space, enabling the identification of new transient candidates via clustering and similarity searches.
2.2 Sparse Representation Learning
We now aim to learn compact and informative representations from the EEttdtdt cubes. To this end, we employ a SAE—an unsupervised neural network model designed to extract structured, low-dimensional features from high-dimensional inputs. By learning to reconstruct the input data while enforcing sparsity in the latent space, the SAE identifies salient patterns and compresses each tensor into a representation that preserves essential temporal and modality-specific features, while remaining insensitive to irrelevant variation (Ranzato et al., 2007; Ng et al., 2011). This includes robustness to noise, systematics, observational artifacts, and context-dependent features introduced by the data acquisition process. Instead, the learned representations are encouraged to capture the most important latent factors that reflect the underlying physical characteristics and meaningful signals encoded in the time series, rather than superficial or incidental aspects of the data.
Formally, let $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{\tau}\times n_{\epsilon}\times n_{\Delta\tau}}$ denote the input tensor for the $i$-th event time series. The autoencoder consists of two parameterized functions: (i) an encoder network $\phi_{\theta}:\mathbb{R}^{n_{\tau}\times n_{\epsilon}\times n_{\Delta\tau}}\rightarrow\mathbb{R}^{d}$ that maps $\boldsymbol{X}_{i}$ to a $d$-dimensional latent vector $\boldsymbol{z}_{i}=\phi_{\theta}(\boldsymbol{X}_{i})$, and (ii) a decoder network $\psi_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n_{\tau}\times n_{\epsilon}\times n_{\Delta\tau}}$ that reconstructs the input as $\hat{\boldsymbol{X}}_{i}=\psi_{\theta}(\boldsymbol{z}_{i})$. To encourage sparsity in the learned representations, we impose an L1 penalty on the latent vector $\boldsymbol{z}_{i}$. The total training loss balances reconstruction accuracy and sparsity, and is given by:
$\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\left\|\boldsymbol{X}_{i}-\hat{\boldsymbol{X}}_{i}\right\|_{2}^{2}+\lambda\left\|\boldsymbol{z}_{i}\right\|_{1},$ (4)
where $\lambda>0$ controls the strength of the sparsity regularization. Once trained, we discard the decoder and use the encoder $\phi_{\theta}$ as a standalone feature extractor. The resulting latent vectors $\boldsymbol{z}_{i}$ serve as learned representations of the original event time series and can be directly used for downstream tasks such as clustering, anomaly detection, or similarity-based retrieval. A similar procedure is followed when using the EEtt maps as inputs to a SAE.
3 Experiments on Downstream Tasks
Application to X-ray Astronomy
We apply our method to data from the Chandra Source Catalog (CSC) version 2.1 (Evans et al., 2024). Specifically, we use region event files, which list individual photon events for a specific source in a given observation—each with associated arrival time and energy. The dataset comprises 95,473 event files from 58,932 unique X-ray sources. For further details on the dataset, please refer to Dillmann et al. 2025. Using the terminology introduced in Section 2.1, the photon arrival times correspond to $\boldsymbol{t}$ and the photon energies to $\boldsymbol{E}$. We choose the event transformation function $f(\cdot)$ to be the base-10 logarithm, i.e., $f(\boldsymbol{E})=\log_{10}(\boldsymbol{E})$, which is commonly used in X-ray astronomy. For the EEttdtdt cubes, we choose a dimensionality of $(n_{\tau},n_{\epsilon},n_{\Delta\tau})=(24,16,16)$. For the SAE, we choose a sparsity strength of $\lambda=0.1$ and its latent vector dimension is 24. We also run the experiments described in this section on the EEtt maps with a dimensionality of $(n_{\tau},n_{\epsilon})=(24,16)$. In this case, we use a SAE with convolution layers and a latent vector dimension of 12. More details are provided in Appendix B.
Figure 2: Nearest-neighbor retrieval results in the learned latent space for three representative transient types: dips from a low-mass X-ray binary (top row, 300 s bins), a flare from a young stellar object (middle row, 400 s bins), and pulsations from a pulsar (bottom row, 100 s bins). For each target light curve (column 1), we show its three nearest neighbors (columns 2–4). The retrieved neighbors correspond to physically similar phenomena: dips from low-mass X-ray binaries, flares from young stars or variable stars, and pulsations from pulsars.
Dimensionality Reduction & Embedding Space
[…]
Semantic Clustering & Unsupervised Classification
By clustering the learned latent space using DBSCAN (Ester et al., 1996), we successfully isolate flare-dominant and dip-dominant groups, as shown in Figure 1C, leading to the identification of 3117 flare and 685 dip candidates. This yields a publicly available catalog of X-ray flares and dips (Dillmann et al., 2025) that required only minimal manual filtering. Both clusters demonstrate high purity, with the flare cluster in particular containing a very high proportion of genuine flaring sources. Notably, we discover a previously unreported extragalactic FXT with unique temporal and spectral properties (Dillmann et al., 2025), and a tidal disruption event (TDE) from a newly identified hyperluminous supersoft X-ray source (Sacchi et al., 2025), both missed by previous transient searches in the Chandra archive (Yang et al., 2019; Quirola-Vásquez et al., 2022).
Anomaly Detection & Similarity Search
Transient sources such as flares and dips represent only a small fraction of the dataset, which is largely dominated by relatively steady sources, making them anomalies from a data science perspective. The learned representations support anomaly detection and similarity-based retrieval of X-ray sources with distinct temporal and spectral features. In Figure 2, we demonstrate this by retrieving the three nearest neighbors using $k$-nearest neighbors (Cover & Hart, 1967) for three target sources: a dip from a low-mass X-ray binary, a flare from a young stellar object, and a pulsar. For each target source, the model retrieves similar light curves from the same or a closely related physical class, indicating that the latent space captures physically meaningful structure across source types.
Supervised Prediction Tasks
We train a vanilla XGBoost classifier and regressor (Chen & Guestrin, 2016) on the learned latent representations to distinguish highly variable X-ray sources (variability index >> 6) from non-variable ones (variability index $\leq$ 6), and to predict the hardness ratio of each source. The performances are summarized in Table 1. The classifier achieves a high variability classification accuracy of 0.97, outperforming the method proposed in Song et al. (2025), and the hardness ratio regressor attains a slightly reduced R2 score of 0.76, demonstrating that the learned representations encode temporal variability well and retain substantial predictive power for spectral hardness.
Table 1: Performance summary of XGBoost models for variability classification and hardness ratio prediction. The models use 100 estimators and default hyperparameters without fine-tuning. An 80-20 train-test split is used for evaluation.
Task Metric Value
Variability Classification Accuracy 0.97
F1-score (non-variable) 0.98
F1-score (variable) 0.72
Hardness Ratio Prediction R2 0.76
MSE 0.03
4 Conclusion
We introduce a unified framework for learning sparse representations from event time series to enable a variety of different downstream tasks. We represent irregular event time series as EEttdtdt cube tensors and employ a SAE to learn sparse embeddings that capture physically meaningful variation while increasing their robustness to noise and other irrelevant features. The method requires minimal manual tuning and scales to large datasets across various domains, and serves as a foundation for subsequent analysis like anomaly detection, similarity search, and unsupervised classification. Applied to archival Chandra X-ray data, the learned representations encode temporal and spectral information and enabled the discovery of a new extragalactic FXT and a TDE from a hyperluminous supersoft X-ray source, both previously missed by other transient search methods. Future work involves optimizing the resolution of the EEttdtdt cubes, the deployment of more expressive SAE architectures and the application of the method to a variety of other scientific datasets.
Software and Data
The data used in this paper was obtained from the publicly available CSC, using their public interfaces (https://cxc.cfa.harvard.edu/csc/). The preprocessed datasets are provided here. All intermediate data products can be produced using the code provided in the GitHub repository https://github.com/StevenDillmann/ml-xraytransients-mnras. The catalog of transient candidates is publicly available on Zenodo and VizieR.
Acknowledgements
This research has made use of data obtained from the Chandra Source Catalog provided by the Chandra X-ray Center (CXC). Juan Rafael Martínez-Galarza acknowledges support to AstroAI. We thank the reviewers for their insightful comments.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and its applications to Astrophysics. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A $E$-$t$ Maps and $E$-$t$-dtdt Cubes Visualizations
Figure 3 displays the light curves, EEtt maps, and EEttdtdt cubes for event files featuring a dip, a flare, and pulsations, respectively.
Figure 3: The top row shows 300 s bin light curves from event files featuring a dip (blue), a flare (red), and pulsations (green), respectively. The middle row shows the corresponding EEtt maps, and the bottom row shows the corresponding EEttdtdt cubes for these event files.
Appendix B SAE Architecture and Training
The architecture of the experimental SAE for the EEttdtdt cube inputs is summarized in Table 2. For the experiments presented in this paper, we flatten the input and use fully-connected layers for computational reasons. We recommend experimenting with more sophisticated SAE architectures to fully leverage the value of the EEttdtdt cubes and the presented SAE representation learning approach. Starting with the original dataset of 95,473 samples, we first apply a 90/10 split to obtain 85,925 samples for training and validation, and 9,548 for testing. We then further split the training and validation subset using an 80/20 ratio, resulting in 68,740 training samples and 17,185 validation samples. The model is trained for up to 200 epochs with a batch size of 1024. Optimization is performed using Adam (Kingma & Ba, 2014) with an initial learning rate of $0.01$. A learning rate scheduler reduces the rate by a factor of 10 if the validation loss plateaus for more than 10 epochs. We also apply early stopping based on the validation loss, which halts training if no improvement is observed for 25 consecutive epochs and restores the best-performing model weights. A similar procedure is followed for the experimental SAE with convolution layers summarized in Table 3 that we used for the EEtt map inputs.
Table 2: Summary of the encoder architecture of the fully-connected autoencoder used to extract informative features from the $E$-$t$-dtdt cubes. Each layer uses Leaky ReLU activation and each standard fully-connected layer is followed by batch normalization with momentum 0.9.
Layer Output Shape
Input (24, 16, 16)
Flatten 6144
Dense 1536
Dense 384
Dense 92
Dense (Bottleneck) 24
Table 3: Summary of the encoder architecture of the convolutional autoencoder used to extract informative features from the $E$-$t$ maps. Each layer uses Leaky ReLU activation and each standard fully-connected layer is followed by batch normalization with momentum 0.9.
Layer Output Shape Filters Kernel Stride
Input (24, 16) - - -
Convolution (24, 16) 32 (3, 3) -
Convolution (12, 8) 32 (2, 2) 2
Convolution (12, 8) 16 (3, 3) -
Convolution (6, 4) 16 (2, 2) 2
Flatten 384 - - -
Dense 192 - - -
Dense 48 - - -
Dense (Bottleneck) 12 - - -
Appendix C Embedding Space from the SAE applied on the EEtt Maps
Figure 4 shows the embedding space obtained from the experimental SAE applied on the EEtt maps.
Figure 4: Two-dimensional t-SNE projection of the learned latent space from the SAE applied on the EEtt maps. Panel A: Points are color-coded by the variability index of the corresponding X-ray sources. Panel B: Points are color-coded by the hard-to-soft X-ray hardness ratio. Panel C: Known dips, flares, and pulsating sources (crosses) distributed across the embedding space.